OpenStudy (cherrilyn):

We have calculated a confidence interval based on a sample of size n = 100. Now we want to get a better estimate with a margin of error that is only one-fourth as large. How large does our new sample need to be?

OpenStudy (cherrilyn):

25 50 200 400 Or 1600

OpenStudy (anonymous):

I think it is 25 100 * 1/4

jimthompson5910 (jim_thompson5910):

Let n = original sample size and n' be the new sample size Now let's use the formula n = sigma*(z/E)^2 and replace E with E/4 and see what happens. n = sigma*(z/E)^2 n' = sigma*(z/(E/4))^2 n' = sigma*((4z)/E)^2 n' = sigma*4^2*(z/E)^2 n' = sigma*16*(z/E)^2 n' = 16*sigma*(z/E)^2 n' = 16*n So n' is 16 times larger than the original sample size. Since the original sample size is n = 100, this means that n' = 16*n n' = 16*100 n' = 1600 So the new sample size needs to be 1600 to make the new margin of error be one-fourth of the original margin of error.